Find the mean of the probability distribution of a random variable X which can take


only the values 2, 4, 5, and 9, given that P(2)=9/20, P(5)=1/20, P(5)=1/5, and


P(9)=3/10.

1
Expert's answer
2022-03-22T17:38:58-0400

"\mu=2*\frac{9}{20}+4*\frac{1}{20}+5*\frac{1}{5}+9*\frac{3}{10}=6.8."